### v1

# HREB-CRF: Hierarchical Reduced-bias EMA for Chinese Named Entity Recognition

Abstract
Incorrect boundary division, complex semantic representation, and differences in pronunciation and meaning often lead to errors in Chinese Named Entity Recognition(CNER). To address these issues, this paper proposes HREB-CRF framework: Hierarchical Reduced-bias EMA with CRF. The proposed method amplifies word boundaries and pools long text gradients through exponentially fixed-bias weighted average of local and global hierarchical attention. Experimental results on the MSRA, Resume, and Weibo datasets show excellent in F1, outperforming the baseline model by 1.1%, 1.6%, and 9.8%. The significant improvement in F1 shows evidences of strong effectiveness and robustness of approach in CNER tasks.
Equal contribution  Corresponding author
Keywords NER  $\cdot$ Hierarchical Attention  $\cdot$ RoBERTa  $\cdot$ Token Classification
1 Introduction
Token classification is one of the key research directions in NLP, where NER assigns entity types to each input token for extracting critical information. An effective NER system not only bolsters tasks such as sentiment analysis, compliance monitoring, and recommendation systems, but also enables deeper text comprehension and facilitates automated processing across diverse contexts. Classic approaches like Lattice-LSTM[1] laid the groundwork for structured Chinese NER, while subsequent methods, including Flat-Transformer[2] and BERT-based solutions[3], have demonstrated notable success.
However, CNER poses unique challenges due to the complexity of the language. 1) Lack of Explicit Word Boundaries: Unlike English, Chinese texts are written without whitespace separation, making entity boundary detection especially difficult. 2) Complex Semantics and Polysemy: The rich morphology, polysemy, and extended meanings of Chinese require precise context modeling to avoid confusion caused by homophones or nested entities. 3) Segmentation Ambiguity: Conventional text features (e.g., radicals, phonetics) often struggle with local segmentation errors and global dependencies, leading to suboptimal performance in real-world scenarios.
To address these issues, Exponential Moving Average (EMA)[4] has garnered attention due to its ability to stabilize model updates and smooth out noisy signals, particularly beneficial for local boundary detection in time-series-like or sequential data. By exponentially decaying historical information, EMA can emphasize recent tokens and reduce the impact of earlier noise, which aligns well with the fine-grained boundary identification crucial in Chinese NER. Nevertheless, a simple EMA mechanism often underestimates long-range dependencies, which are vital for handling polysemy and nested structures in Chinese.
Therefore, inspired by the MEGA structure[5], a Reduced-biased Hierarchical EMAttention framework (RHEMA) is proposed integrated with RoBERTa[6]. This hierarchical design, shown as Figure-1 ensures both short-range (local) and long-range (global) dependencies are jointly modeled, addressing segmentation ambiguities. The reduced-bias connection further mitigates the risk of gradient vanishing or exploding when stacking deep layers, thus providing more stable parameter updates.
Figure 1: Constructing on the uniqueness of the decay weight mechanism in the exponential averaging method, a reduced-biased exponential averaging layer is incorporated into the attention mechanism.
To sum up, proposed contributions are mainly as follows:
  1. 1.
    The proposed work innovatively introduces RHEMA into the NER task, enabling more accurate modeling of short-range dependencies for entity boundary detection while preserving global contextual insights.
  2. 2.
    By leveraging an advanced text embedding approach, the model achieves superior efficiency, stability, and generalization capabilities in text vectorization and complex language representation tasks, thereby significantly enhancing overall performance.
  3. 3.
    The proposed work apply a reduced-bias method in each module to alleviate gradient vanishing, ensuring stable gradient flow and further enhancing the model’s ability to capture intricate linguistic features in deep architectures.
2 Related Work
Deep learning methods have significant advantages in named entity recognition tasks. One of its main advantages is the ability to capture contextual information by learning nested representations, helping to identify the boundaries of entities more accurately. Deep learning models can automatically learn abstract features in text without manually designing features, thereby more comprehensively understanding the multi-faceted information of entities, including morphology, syntax, and semantics. In addition, deep learning models can handle long-distance dependencies, capture dependencies at different distances in text through operations such as recursion or convolution, and improve the understanding of the entire text structure, thus improving the accuracy of named entity recognition. And it can also be combined with conditional random fields (CRF) to globally model the label sequence, consider the joint probability of the overall sequence, ensure the consistency between labels, and avoid unreasonable label sequences. This integration method makes the model better in NER. More comprehensive and accurate identification of entities in tasks.
Recently development of these methods shows a more comprehensive understanding of contextual information and improves the generalization ability in multiple fields and multiple contexts.
Qiu et al. [7] highlighted that discrepancies in text semantics caused by inaccurate tokenization are a distinctive challenge in Chinese named entity recognition (NER) compared to English NER. Different tokenization schemes can lead to varying evaluation metrics. He et al.’s study [8] found that character-based segmentation in Chinese NER outperforms word-based segmentation, indicating that a finer granularity is advantageous for Chinese language processing. Further notable findings show that static embedding techniques, such as Word2Vec [9], while effective in deep neural networks, fail to address polysemy—a characteristic unique to Chinese language processing. Building on this, the CWVF-BiLSTM-CRF model developed by Ye et al. [10] demonstrates that incorporating both character and word inputs can reduce accuracy errors caused by tokenization while improving entity recognition precision.
Cao et al. [11] uses an adversarial transfer learning model to predict text boundary intervals, which effectively filters noise within sentences. Lattice-LSTM [1] is a pioneering method in Chinese NER, as prior research consistently shows that dictionary-based approaches can significantly enhance Chinese NER performance, with many studies adapting their models based on this algorithm. To address lexical conflicts, Gui et al. [12] introduced the LR-CNN model, which employs CNN layers for feature extraction and integrates attention mechanisms between CNN layers to manage vocabulary conflicts through a "Rethinking" mechanism. Additionally, Liu’s work [13] encodes word information using multiple representations, converting words into fixed-length feature vectors. Lastly, the FLAT-Transformer model by Li et al. [2] leverages a lattice-based structure to represent text, converting character grids into a flattened representation that enhances NER performance.
Moreover, specialized data structures can effectively capture named entity features. The Tree-LSTM model [14] demonstrates that using a topological network structure within data structures can not only integrate sub-feature information but also provide richer topological representations. DyLex [15], proposed by Wang et al., is a BERT-based dictionary augmentation method that addresses noise in text sequences under supervised training and allows dictionary updates without requiring retraining. Liu [16] introduced the TFM network, aimed at enhancing character understanding by integrating positional information, thereby capturing both relative and absolute positional relationships among characters within the model to improve character interrelation comprehension.
Besides, many studies also explore Chinese word structure, phonetics, and radicals as features for named entity recognition. Mai [17] examines different practical features of homophones based on lexical and phonetic characteristics. Li [18] leverages syntactic dependency trees to guide global attention mechanisms through rule-based methods.
3 Methodology
3.1 NER Task Definition
The NER task is a typical sequence labeling task. Based on a $token$ sequence $x={x_{1},x_{2},\cdots,x_{n}}$ of a given length $n$, the output $y={y_{1},y_{2},\cdots,y_{n}}$ is predicted. Where $x_{i}\in\text{T}$, $T$ represents the vocabulary. In this study, the BIO labeling method is used to label the category $y$. The entire task flow requires the construction of a text processing sequence, as shown in the Equation-1.
$input$ =$=\{x_{1},x_{2},\cdots,x_{n}\}$ Tokenization (1)
$vectoried$ =$=\{\text{i}_{1},\text{i}_{2},\cdots,\text{i}_{n}\}$ BERT Embedding
$featured$ =$=\{\text{v}_{1},\text{v}_{2},\cdots,\text{v}_{n}\}$ Feature Extraction
$cls$ =$=\{\text{f}^{1}_{1},\text{f}^{1}_{2},\cdots,\text{f}^{C}_{n}\}$ Segmentation Weight
$y$ =$=\{c_{1},c_{2},\cdots,c_{n}\}$ Sequence Dependency
where $C$ denotes number of classes.
In detail, the whole process starts with the input text sequence, and the text is converted into a set of $token$ through word segmentation. Then, these $token$ are mapped to the vector space through the BERT embedding layer to generate an embedded representation of each $token$. Subsequently, the embedded representation is passed to the feature extraction module to obtain deeper contextual features. The classification layer calculates multi-category scores for the feature vectors at each position to obtain the weights corresponding to each category. Finally, CRF layer is used to further model the dependencies between sequences, and the argmax operation is used to select the label sequence with the highest probability to obtain the final prediction output.
3.2 Overview of HREB-CRF Architecture
Proposed model is shown in Figure-2. BERT structure model is used as the pre-training head, and output the start and end (B-Cls and last I-Cls structure annotation) features of the CRF-constrained entity. On this basis, the additional Hierarchical EMA module is used to enhance capturing the short-distance relationship of the text vector. Compared with traditional attention method, this improved method has the ability to extract both short-range and long-range features. Next, this hierarchical dependency weights are input into the Bi-LSTM layer to capture the context and strengthen the long-distance dependency capabilities, and finally input into the CRF for sequence dependency decoding. In addition, a reduced-biased module is added to each layer to reduce the possibility of gradient disappearance.
Figure 2: Overall Structure of proposed model. The left figure shows the end-to-end model architecture from text input to entity classification output, and the right figure shows the data processing flow of the RHEMA architecture.
3.3 Exponentially Moving Average
Moving average method is widely used in sequence analysis tasks, such as stock and anomaly detection. It is proposed to reduce the global impact of local outliers anomalies. Generally speaking, most of the usage of MA are: simple moving average(Equation-2), weighted moving average(Equation-3), cumulative moving average(Equation-4) and exponentially weighted moving average(Equation-5).
$SMA_{t}=\frac{1}{N}\sum_{i=t-N+1}^{t}x_{i}$ (2)
$WMA_{t}=\frac{\sum_{i=0}^{N-1}w_{i}\odot x_{t-i}}{\sum_{i=0}^{N-1}w_{i}}$ (3)
$CMA_{t}=\frac{\sum_{i=1}^{t}x_{i}}{t}$ (4)
$EMA_{t}=\alpha\odot x_{t}+(1-\alpha)\odot EMA_{t-1}$ (5)
In the exponentially moving average (EM) formula, $\alpha$ is used as a weighting factor to determine the degree of attenuation of historical data. A larger $\alpha$ value indicates that current data is given more importance and the reliance on earlier data is weakened. Therefore, EM can effectively reduce the impact of random factors in the sequence on the overall trend, because as time goes by, $(1-\alpha)^{t}\cdot y_{t}$ will gradually decrease and tend to zero, gradually diluting the weight of earlier data.
The advantage of EM lies in its high responsiveness to recent data, which can adapt to data trend changes more quickly while reducing lag. This feature is particularly suitable for capturing turning points and mutations in trends in sequences, especially for application scenarios such as time series data with turning points that require rapid weight adjustment. In the NER task, finding entity boundaries is one of the key issues. In addition, EM reduces the impact of long-term historical data through exponential decay, and can smooth noise without losing data continuity, making it robust in dynamic environments.
Similar to CNNs, EM can be regarded as applying additional decaying weights to each data block. In CNNs, the convolution kernel extracts local features through a sliding window, while in EMA, the weighting factor $\alpha$ plays a similar role to the convolution kernel, giving different importance to data of different time steps through gradually decaying weights. This decaying weight mechanism makes EM pay more attention to recent data and reduces the impact of early data on the overall mean. It is similar to the weight sharing and local perception characteristics in convolution operations. This feature gives EM sensitivity to changes in data trends, can smooth short-term fluctuations in the sequence, and retain information about longer trends, thereby playing a role in smoothing and denoising in the analysis and prediction of time series data.
3.4 HEMA
Since the exponential averaging method is unique in the decay weight mechanism, we add the reduced-biased exponential averaging layer to the attention mechanism. Based on the MEGA method of [5], reduced-biased Hierarchical EMA is introduced as improved model. This improved architecture is shown in the Figure-3.
Figure 3: Proposed RHEMA architecture with added reduced-biased module. The figure illustrates the architecture of the RHEMA framework for text processing, enhanced by the addition of a reduced-bias module, marked in red. The input text is initially transformed through a token embedding layer. The core RHEMA block includes sequential processing through Batch Normalization and a FeedForward network. A novel "Reduced-Bias Add" module, shown with red arrows, is integrated to mitigate potential bias in the output, while preserving a feedback loop to earlier stages. This enhancement aims to improve robustness and fairness in predictions. The final output weights are generated post these computations.
RHEMA uses GAU[19] to share feature dimensions. First, pass it to the EMA layer at Equation-6, then use the SiLU self-gated activation function to linearly project to z dimension, and add nonlinear features.
$X_{EMA}=EMA(X_{in})$ (6)
$Z=F_{SiLU}(X_{EMA}W_{z}+b_{z})+X_{in}$ (7)
With the advantage of SiLU, the operation here can compress the originally redundant d-dim input into a more compact z-dim representation while retaining the context information. Equation-8 represents the specific calculation process of SiLU.
$F_{SiLU}=\sigma(x)+x\cdot\sigma(x)\cdot(1-\sigma(x))$ (8)
Then, through the output of Equation-7, Query(Equation-9), Key(Equation-10) and Value(Equation-11) is calculated. Among them, the bias terms ($\mu_{q}$, $\mu_{k}$ and $b_{v}$) add additional degrees of freedom to the attention mechanism, which allows the query and key to adjust their respective baseline values based on the shared representation $Z$ to make them more suitable for broader tasks. And $v$ refers to the extension dim of value series.
$Q=\mathcal{K}_{q}\odot Z+\mu_{q}$ (9)
$K=\mathcal{K}_{k}\odot Z+\mu_{k}$ (10)
$V=F_{SiLU}(X_{in}W_{b}+b_{v})$ (11)
Next, the result of the key-value pair query can be expressed as Equation-12.
$O=f_{S}(\frac{QK^{T}}{\daleth}+b_{rel})V$ (12)
where $f_{S}(x)$ means single headed attention, and the bias term of $b_{rel}$ can be optimally solved by attention. $\daleth$ represents the scaling factor, which reduces the digits of large queries.
by simulating the gating mechanism of GRU, activation output(Equation-13) is introduced. It affects the influence of each cell on the global affection.
$\hat{Y}=F_{SiLU}(X_{SiLU}W_{h}+(\gamma\odot O)U_{h}+b_{h})$ (13)
Where the reset gate $\gamma$ controls whether the weight matrix $U_{h}$ affects the final output.
Finally, the output is represented as Equation-14.
$Y=\Phi\odot\hat{Y_{s}}+(1-\Phi)\cdot X_{in}$ (14)
where $\Phi$ represents the update gate, smoothing the result by Sigmoid activation function.
The single-head attention calculation method mentioned in Equation-14 can be expressed as Figure-4.
Figure 4: The architecture of single-head attention.
The output $Y$ combines the candidate activation state $\hat{Y}$ and the original input $X_{in}$, and the update gate $\Phi$ controls the information fusion ratio between the two parameters. When calculating the candidate activation state $\hat{H}$, the exponential average attention mechanism uses the Laplace attention function(Equation-15) as the core component. Based on the concept of reduced-bias, it can be improved the reduced-bias Laplace function as Equation-16 to adapt to the introduction of deviation inside.
$F_{la}(x;\mu,\sigma)=\frac{1+\text{erf}\left(\frac{x-\mu}{\sigma\sqrt{2}}\right)}{2}$ (15)
$F_{rla}(x;\mu,\sigma)=Norm(\frac{1+\text{erf}\left(\frac{x-\mu}{\sigma\sqrt{2}}\right)}{2}+X_{in})$ (16)
The Laplace attention function maps the attention score of $QK^{T}$ to a smooth range by dynamically adjusting the parameters $\mu$ and $\sigma$, providing a more stable gradient and a more flexible attention distribution. This smoothing range is mapped to (0, 1) using the erf function, which improves the global context modeling ability of the model, and ultimately improves the expression effect and task performance of $Y$.
3.5 Reduced-bias Module
Compared with the traditional residual connection layer, proposed work have improved on the basis of the classic residual network structure (Residual Network)[20], and added a dynamic adjustment mechanism for the network weights before and after the residual connection.
Specifically, the proposed Reduced-bias module fits the sensitivity and influence of different weight parameters on downstream gradient updates by constructing an adaptive weight adjustment strategy during the forward propagation process. In the verification feedback stage, the module can select the optimal parameter configuration through dynamic optimization, thereby enhancing the gradient transfer efficiency and learning ability of the model. In addition, this improvement can effectively alleviate the information transfer bias problem caused by fixed weights in traditional residual connections, and provide stronger robustness and flexibility for deep feature extraction and learning of the model.
Classic resnet can be expressed as Equation-17.
$\mathbf{y}_{l}=\mathcal{F}(\mathbf{x}_{l},\mathbf{W}_{l})+\mathbf{x}_{l}$ (17)
where $\mathbf{x}_{l}$ is the input of the $l$-th layer module. $\mathcal{F}(\cdot)$ represents the operation of the network in this layer. $\mathbf{W}_{l}$ is a parameter of the $l$-th module.
Among Equation-17, dynamic adaptation mechanism parameters $\alpha$ and $\beta$ is introduced. The new network can be expressed as Equation-18.
$\mathbf{y}_{l}=\alpha_{l}\cdot\mathcal{F}(\mathbf{x}_{l},\mathbf{W}_{l})+\beta_{l}\cdot\mathbf{x}_{l}$ (18)
where $\alpha$ controls the influence of the residual term, and $\beta$ controls the influence of the input direct connection term. Besides, $l\in L$, $L$ represents all the layer connected by Reduced-biased module.
$t_{l}=\sigma\left(\frac{\partial\mathcal{L}}{\partial\mathcal{F}(\mathbf{x}_{l},\mathbf{W}_{l})}\cdot\mathbf{W}_{t}+b_{t}\right)$ (19)
where $\mathcal{L}$ is the loss function, $\mathbf{W}_{\alpha}$ and $\mathbf{W}_{\beta}$ are learnable weight matrices, $b_{\alpha}$ and $b_{\beta}$ are bias terms, and $\sigma(\cdot)$ is the activation function.
By deriving Equation-19, $\mathbf{y}_{l}$ can be got by Equation-20.
$\mathbf{y}_{l}=$ $\,\sigma\left(\frac{\partial\mathcal{L}}{\partial\mathcal{F}(\mathbf{x}_{l},\mathbf{W}_{l})}\cdot\mathbf{W}_{\alpha}+b_{\alpha}\right)\cdot\mathcal{F}(\mathbf{x}_{l},\mathbf{W}_{l})$ (20)
$+\sigma\left(\frac{\partial\mathcal{L}}{\partial\mathbf{x}_{l}}\cdot\mathbf{W}_{\beta}+b_{\beta}\right)\cdot\mathbf{x}_{l}$
Using optimization of the parameters of this layer, the contribution of different paths in the residual connection to the final output can be dynamically adjusted, thus achieving adaptive learning and weight optimization.
3.6 Robustly Optimized BERT
RoBERTa (Robustly Optimized BERT Approach)[6] is an improved version of BERT proposed by Facebook AI. It greatly improves the performance of the model by optimizing the training strategy and hyperparameters. Compared with BERT, RoBERTa has been enhanced and improved in many aspects. First of all, RoBERTa uses a larger data set for training, and the total amount of data is more than ten times larger than BERT. This significantly improves the language representation capabilities of the model.
In addition, RoBERTa removes the Next Sentence Prediction task in BERT and only retains the training target based on Masked Language Model. Research shows that removing NSP does not reduce model performance, but reduces unnecessary constraints, thereby improving task performance. At the same time, RoBERTa introduces a dynamic masking strategy instead of BERT’s static masking. This means that the mask is re-randomly generated for the input text each time, thereby enhancing the diversity of training and the robustness of the model.
Through the above improvements, RoBERTa has achieved excellent performance on multiple natural language understanding tasks, including benchmarks such as GLUE, RACE, and SQuAD. Compared with BERT, RoBERTa’s more powerful expression ability and flexible training strategy make it a pre-trained language model with outstanding performance in many downstream tasks.
4 Experiments
4.1 Datasets Description
Three NER benchmark datasets are collected to verify the performance of proposed model. They are from Weibo[21], MSRA[22], and Resume[1].
The specific properties of the original datasets are shown in the Table-1. It should be noted that since MSRA does not have a validation set by default, all experiments directly use the test-set as the valid-set. For the other two datasets, a split rate of 7:1.5:1.5 is used to evaluate.
Table 1: Properties of the three datasets used.
datasets weibo msra resume
class 8 3 8
train 1350 46364 3821
test 270 4365 477
valid 269 4365 463
avg length 54.61 46.80 32.47
max length 175 581 178
min length 7 5 3
4.2 Experiment Settings
To verify the versatility and effectiveness of the proposed model in the NER task, BiLSTM-CRF is used BiLSTM-CRF as the baseline model. On this basis, comparative experiments and ablation experiments are designed. In addition, it is also compared the impact of different pre-training heads on model indicators.
Experimental environment is based on the Ubuntu 22.04LTS operating system, using the NVIDIA RTX4090 for inference and computing, and the cuda version is 12.4, with Python 3.9 environment. During the training process, Adam is used to optimize the learning rate. The number of training epochs is set to 100 times with early stop callback.
4.3 Evaluation Metrics
Proposed work uses precision (P), Recall (R), and F-Score ($F_{1}$) to evaluate the model.
The negative log-likelihood function, shown as Equation-21 is considered as the loss function to quantitatively evaluate the performance of the model.
$\text{NLL}=-\sum_{i=1}^{N}\sum_{c=1}^{C}y_{i,c}\log(\hat{y}_{i,c})$ (21)
where $y_{i,c}$ is the binary indicator (0 or 1) of whether class label $c$ is the correct classification for sample $i$, and $\hat{y}_{i,c}$ is the predicted probability of sample $i$ belonging to class $c$.
4.4 Experimental Results
4.4.1 Comparison Studies
The experimental results on MSRA, Resume and Weibo for comparison studies are shown in the several tables respectively.
Table 2: Comparison Experimental results on MSRA.
Models Description MSRA
P R F1
BiLSTM+CRF 84.65 80.68 82.62
Lattice LSTM Zhang 2018[1] 93.57 92.79 93.18
AS Cao 2018[11] 91.73 89.58 90.64
TextCNN Jia 2018[23] 91.63 90.56 91.09
DEM Zhang 2019[24] 90.59 91.15 90.87
MFE Li 2021[25] 90.47 84.70 87.49
MSFM Zhang 2022[26] 86.90 85.46 86.17
DAE Liu 2024[27] 92.73 90.15 91.42
HREB (Ours) 94.58 93.98 94.28
Compared with multiple existing methods on MSRA dataset by Table-2, he classic Lattice LSTM model achieves an F1 score of 93.18, demonstrating its effectiveness by combining character and word-level information. However, proposed method surpasses this benchmark with an F1 score of 94.28, demonstrating stronger semantic modeling capabilities. Compared with the 82.62 F1 score of BiLSTM+CRF as the baseline model, proposed structure shows significant advantages in deep feature extraction and context information utilization.
In addition, compared with DEM and DAE in recent years, proposed model has significantly improved both precision and recall, especially showing stronger generalization ability in complex entity recognition tasks. Comprehensive results show that proposed method achieves state-of-the-art performance on MSRA dataset.
Table 3: Comparison Experimental results at Resume.
Models Description Resume
P R F1
BiLSTM+CRF 94.79 93.68 94.23
Lattice LSTM Zhang 2018 [1] 94.81 94.11 94.46
LLPA Gan 2021 [28] 96.36 94.88 95.62
MSFM Zhang 2022 [26] 96.08 94.79 95.43
MFT Han 2022 [29] 96.05 95.52 95.78
DAE Liu 2024 [27] 96.92 95.18 96.04
HREB (Ours) 95.53 96.61 96.07
Although proposed model has achieved meaningful recall (96.61) and F1 score (96.07) on the Resume dataset, detailed in Table-3. There are still some shortcomings that deserve attention. First, compared with the DAE model, the precision is slightly lower (95.53 vs. 96.92), indicating that there is space for improvement in the prediction accuracy of specific entities.
Table 4: Comparison Experimental results at Weibo.
Models Description Weibo
P R F1
BiLSTM+CRF 57.84 49.08 53.10
Lattice LSTM Zhang 2018[1] 53.04 62.25 58.79
Glyce Meng 2019[30] 67.68 67.71 67.60
FLAT Li 2020[2] 63.42
FLAT+BERT Li 2020[2] 68.55
MSFM Zhang 2022[26] 60.75 51.83 55.94
DAE Liu 2024[27] 69.68 48.89 57.45
HREB (Ours) 66.57 70.76 68.60
Figure 5: Based on the web search results, a piece of news was randomly selected as a test sample, which contains the extraction of names of people, places and organizations.
Proposed method achieves an F1 score of 68.60, slightly higher than FLAT+BERT (68.55) and close to Glyce (67.60), demonstrating robustness in complex text processing on Weibo, shown as Table-4. Furthermore, proposed method performs well in recall (70.76), indicating its advantage in identifying more potential entities. However, compared with the precision of DAE (69.68), proposed method still has a certain gap in accuracy, which suggests that proposed model still has room for improvement in accurate prediction when dealing with noisy social media data.
4.4.2 Ablation Studies
In the ablation experiment, shown as Table-5, two experiments are designed to evaluate the real impacts of proposed structure. First, each module is removed to confirm the synergy between different modules. Second, BERT is replaced or removed to compare the impact of word embedding weights on the next layer input.
Comparison between the base model and different BERT models
The experimental results of comparing the base model (S.N.1) with the models that introduce different BERT variants (S.N.2 and S.N.3) show that RoBERTa outperforms BERT in terms of vocabulary vectorization. This shows that RoBERTa has a stronger ability to represent text semantics and can capture contextual information more effectively.
Comparison between naive attention and hierarchical EMA
When comparing the naive attention model (S.N.4) with the model based on hierarchical exponential moving average (Hierarchical EMA) (S.N.5), the experimental results show that the hierarchical attention mechanism can more effectively extract the weight distribution of key information under the constraint of exponential gain weights.
Comparison between static and dynamic reduced-bias models
By comparing the immutable Reduced-bias model (S.N.4) with the self-learning Reduced-bias model (S.N.5), the results verify that the dynamically changing coefficients can be adaptively adjusted according to the input text features, thereby improving the model’s ability to express features.
Comparison of the proposed model with the naive BERT model
In the comparison of the final model (S.N.8) with the ordinary BERT model (S.N.7), the experimental results show that even with the introduction of a rich feature extraction module, RoBERTa can still efficiently capture the deep features of the text and show a superior feature learning ability.
Table 5: Ablation Studies on MSRA
S.N. Module Metric
A H R BERT P R F1
1 84.65 80.68 82.62
2 BERT-base 91.02 92.09 91.55
3 RoBERTa 92.21 92.57 92.38
4 $\surd$ 87.42 85.78 86.60
5 $\surd$ 91.35 90.62 90.98
6 $\surd$ $\surd$ 92.92 93.09 93.01
7 $\surd$ $\surd$ BERT-base 93.55 93.88 93.71
8
$\surd$ $\surd$ RoBERTa 94.58 93.98 94.28
4.5 Parameter Selection
In the design of the attention module based on the EMA (Exponential Moving Average) mechanism, the number of attention heads is considered to be one of the important factors affecting the model performance. In order to determine the optimal configuration of attention heads, different numbers of attention heads (16, 32, 64, 128) is set in the experiment and evaluated the performance of the model on the validation set. The experimental results show that when the number of attention heads is set to 64, the model achieves the best performance in performance indicators such as precision, recall, and F1 score.
Table 6: Effects of different parameters on performance in RHEMA.
EMA head metric dataset
MSRA Weibo Resume
n_head 8 P 92.42 65.96 92.24
R 93.44 68.24 94.96
F1 92.93 67.08 93.58
16 P 93.73 65.90 94.08
R 94.59 71.70 95.31
F1 94.16 68.67 94.69
32 P 93.89 66.27 94.52
R 94.86 70.44 95.92
F1 94.37 68.29 95.21
64 P 94.58 66.57 95.53
R 93.98 70.76 96.61
F1 94.28 68.60 95.07
128 P 93.92 64.77 95.16
R 94.50 71.70 95.74
F1 94.21 68.06 95.45
As seen from the Table-6, too few attention heads with 16 or 32, may cause the model to be unable to fully capture feature representations, thereby limiting the performance of the model. When the number of attention heads increases to 64, the performance of the model reaches its peak. However, when the number of attention heads is further increased to 128), the performance of the model does not continue to improve, but instead shows a slight decline. This phenomenon may be due to the fact that too many attention heads increase the complexity of the model, thereby introducing the risk of overfitting.
4.6 Case Analysis
Real Example
Select text data on the Internet for analysis.
In Example-1 of Figure-5, the performance of the model trained on MSRA dataset is verified using normal press releases. The results show that proposed model can effectively extract entity information for regular text. For the baseline model, it recognizes the text of the school-college combination as two different entities. When considering ground truth, it is generally believed that school+college is a complete entity.
Figure 6: Cases in the MSRA validation set are selected for analysis.
Figure 7: In the classic Chinese story “The Legend of the White Snake”, Bai Suzhen and Xiaoqing are female protagonists who were transformed from poisonous snakes into human form. This case involves the extraction of obfuscated names.
In Example-2 of Figure-7, an extreme scenario was constructed. The same Chinese character was expressed as a noun, a verb, and an adjective simultaneously. When these were combined into a sentence, it was verified whether the model could output correctly in extreme cases. The experimental results indicate that the two protagonists could be effectively identified by the proposed model. However, for the baseline model, although the two protagonists were covered, the entity boundary was found to be unclear.
Dataset Example
Select text data on the datasets for analysis.
The analysis presented in Figure-6 addresses two primary challenges in entity recognition, as outlined in the introduction: 1) ambiguity in boundary identification (example-3); 2) errors in entity recognition (example-4). A comparative evaluation demonstrates that the proposed model effectively resolves these issues in the selected examples, exhibiting superior performance relative to the baseline model.
4.7 Analysis and Discussion
Through comparative experiments and ablation studies, an in-depth analysis of various aspects of model performance was conducted. The following are the main findings and discussions derived from the experimental results:
1) Proposed method can be matched with the comparative experiment and the ablation experiment to prove that the model can adapt to different types of semantic information under different scene tasks and adaptively improve the robustness of CNER. Comparing the experimental results of Table-2, Table-4, and Table-3, proposed model achieves the best performance in both regression and F1 indicators. The ablation experiment in Table-5 shows that the hierarchical exponential attention and reduced-biased modules can effectively expand the deep dimensions of the data compared with the baseline model. It performs well in the conventional tasks in Figure-5.
2) Proposed research shows that the Hierarchical Reduced-bias EMA architecture can bring accurate feature extraction for CNER tasks. The modeling of the attention mechanism is to understand the text content from the vector dimension, which is used in the latest LLMs models (such as GPT, Llama). In the field of transfer learning, inspired by DAE[27], proposed model can be used in other sequence modeling tasks. The sequence modeling task uses BERT to model the text vectorization input, and the output is the probability weight of the same length as the input text. It is believed that the model can be applied to NER tasks such as text understanding, text classification, and emotion recognition simultaneously.
3) Increasing the number of attention heads will significantly increase the number of parameters in the multi-head attention mechanism, thereby increasing the training time and inference time of the model. When selecting the number of attention heads, proposed study comprehensively considered the trade-off between performance and efficiency and selected EMA head of 64.
5 Conclusion
Proposed study focus on CNER model that combines RoBERTa, reduced-biased Hierarchical EMA, BiLSTM, and CRF, which conducts extensive experimental verification on multiple Chinese datasets. Experimental results show that proposed method has achieved leading performance on MSRA, Resume, and Weibo datasets, demonstrating strong adaptability to diverse language scenarios. Despite the excellent experimental results, there are still some directions worth exploring. In the future, stronger feature fusion strategies, more robust pre-trained models, and regularization techniques that adapt to diverse scenarios will be explored to further improve model performance and generalization capabilities.
Acknowledgments
This was was supported in part by Agency for Science, Technology and Research.
Appendix
5.1 CRF Decoding
CRF was introduced to strengthen the shortcomings of HMM in feature selection and dependency assumptions. HMM is a generative model that models the joint probability $P(x,y)$, that is, the joint distribution of the observed data and the label sequence, expressed as the Equation-22. This requires HMM to make strong assumptions about the relationship between the observed data and the hidden state (such as the independent and identically distributed assumption and the Markov assumption). In contrast, CRF is a discriminative model that directly models the conditional probability $P(y|x)$ without making independence assumptions on the observed data. Therefore, CRF introduces feature functions and weights into the model and maximizes the conditional probability of given observed data by learning these parameters. It is able to directly model the label sequence without assuming the Markov property between the observed data and the hidden state. This enables CRF to better capture long-distance dependencies in the sequence and automatically learn the relationship between features without manually designing features.
$P(x,y)=P(y_{1})\prod_{t=2}^{T}P(y_{t}|y_{t-1})\prod_{t=1}^{T}P(x_{t}|y_{t})$ (22)
The calculation process of CRF involves constructing and optimizing the conditional probability $P(y|x)$, that is, the probability of predicting the label sequence $y$ given the observation sequence $x$. The core of CRF is to maximize this conditional probability by defining feature functions and learning weights. Specifically, the derivation formula of CRF is:
$P(y|x)=\frac{1}{Z(x)}\exp\left(\sum_{t=1}^{T}\sum_{k}\lambda_{k}f_{k}(y_{t-1},y_{t},x,t)\right)$ (23)
where $f_{k}(y_{t-1},y_{t},x,t)$ is the feature function that defines the relationship between the label sequence $y$ and the observation sequence $x$. $\lambda_{k}$ is the weight parameter of the feature function, which is learned through training. $Z(X)$ is the normalization term (partition function) that ensures that the sum of the conditional probabilities is the Equation-24.
$Z(x)=\sum_{y}\exp\left(\sum_{t=1}^{T}\sum_{k}\lambda_{k}f_{k}(y_{t-1},y_{t},x,t)\right)$ (24)
5.2 Model Training Performance
Figure-8, Figure-9 and Figure-10 are the training processes of MSRA, Weibo and Resume respectively. The training process shows that MSRA reaches asymptotic performance at about 60 steps, Weibo reaches asymptotic performance at 15 steps, and Resume reaches asymptotic performance at 85 steps.
Figure 8: Training progress metrics for MSRA.
Figure 9: Training progress metrics for Weibo.
Figure 10: Training progress metrics for Resume.

### v2

@@
 The significant improvement in F1 shows evidences of strong effectiveness and robustness of approach in CNER tasks.
-Equal contribution  Corresponding author
-Keywords NER  $\cdot$ Hierarchical Attention  $\cdot$ RoBERTa  $\cdot$ Token Classification
-1 Introduction
+Code available at https://github.com/StanleySun233/HREB-CRF.
+Index Terms:
+NER, Hierarchical Attention, RoBERTa, Token Classification
+I Introduction
 Token classification is one of the key research directions in NLP, where NER assigns entity types to each input token for extracting critical information.
@@
 Therefore, inspired by the MEGA structure[5], a Reduced-biased Hierarchical EMAttention framework (RHEMA) is proposed integrated with RoBERTa[6].
-This hierarchical design, shown as Figure-1 ensures both short-range (local) and long-range (global) dependencies are jointly modeled, addressing segmentation ambiguities.
+This hierarchical design, shown as Figure 1 ensures both short-range (local) and long-range (global) dependencies are jointly modeled, addressing segmentation ambiguities.
 The reduced-bias connection further mitigates the risk of gradient vanishing or exploding when stacking deep layers, thus providing more stable parameter updates.
-Figure 1: Constructing on the uniqueness of the decay weight mechanism in the exponential averaging method, a reduced-biased exponential averaging layer is incorporated into the attention mechanism.
-To sum up, proposed contributions are mainly as follows:
+The contributions of this work can be mainly summarized as the follows:
   1. 1.
@@
   2. 2.
+    The proposed work apply a reduced-bias method in each module to alleviate gradient vanishing, ensuring stable gradient flow and further enhancing the model’s ability to capture intricate linguistic features in deep architectures.
+  3. 3.
     By leveraging an advanced text embedding approach, the model achieves superior efficiency, stability, and generalization capabilities in text vectorization and complex language representation tasks, thereby significantly enhancing overall performance.
-  3. 3.
-    The proposed work apply a reduced-bias method in each module to alleviate gradient vanishing, ensuring stable gradient flow and further enhancing the model’s ability to capture intricate linguistic features in deep architectures.
-2 Related Work
-Deep learning methods have significant advantages in named entity recognition tasks.
-One of its main advantages is the ability to capture contextual information by learning nested representations, helping to identify the boundaries of entities more accurately.
-Deep learning models can automatically learn abstract features in text without manually designing features, thereby more comprehensively understanding the multi-faceted information of entities, including morphology, syntax, and semantics.
-In addition, deep learning models can handle long-distance dependencies, capture dependencies at different distances in text through operations such as recursion or convolution, and improve the understanding of the entire text structure, thus improving the accuracy of named entity recognition.
-And it can also be combined with conditional random fields (CRF) to globally model the label sequence, consider the joint probability of the overall sequence, ensure the consistency between labels, and avoid unreasonable label sequences.
-This integration method makes the model better in NER.
-More comprehensive and accurate identification of entities in tasks.
+II Related Work
+Deep learning performs well in NER by capturing contextual information, learning abstract features automatically, and handling long-distance dependencies through recursive or convolutional operations.
+Integrating conditional random fields (CRF) further enhances performance by modeling label sequences globally, ensuring consistency and improving accuracy.
 Recently development of these methods shows a more comprehensive understanding of contextual information and improves the generalization ability in multiple fields and multiple contexts.
-Qiu et al.
-[7] highlighted that discrepancies in text semantics caused by inaccurate tokenization are a distinctive challenge in Chinese named entity recognition (NER) compared to English NER.
-Different tokenization schemes can lead to varying evaluation metrics.
-He et al.’s study [8] found that character-based segmentation in Chinese NER outperforms word-based segmentation, indicating that a finer granularity is advantageous for Chinese language processing.
-Further notable findings show that static embedding techniques, such as Word2Vec [9], while effective in deep neural networks, fail to address polysemy—a characteristic unique to Chinese language processing.
-Building on this, the CWVF-BiLSTM-CRF model developed by Ye et al.
-[10] demonstrates that incorporating both character and word inputs can reduce accuracy errors caused by tokenization while improving entity recognition precision.
-Cao et al.
-[11] uses an adversarial transfer learning model to predict text boundary intervals, which effectively filters noise within sentences.
-Lattice-LSTM [1] is a pioneering method in Chinese NER, as prior research consistently shows that dictionary-based approaches can significantly enhance Chinese NER performance, with many studies adapting their models based on this algorithm.
-To address lexical conflicts, Gui et al.
-[12] introduced the LR-CNN model, which employs CNN layers for feature extraction and integrates attention mechanisms between CNN layers to manage vocabulary conflicts through a "Rethinking" mechanism.
-Additionally, Liu’s work [13] encodes word information using multiple representations, converting words into fixed-length feature vectors.
-Lastly, the FLAT-Transformer model by Li et al.
-[2] leverages a lattice-based structure to represent text, converting character grids into a flattened representation that enhances NER performance.
-Moreover, specialized data structures can effectively capture named entity features.
-The Tree-LSTM model [14] demonstrates that using a topological network structure within data structures can not only integrate sub-feature information but also provide richer topological representations.
-DyLex [15], proposed by Wang et al., is a BERT-based dictionary augmentation method that addresses noise in text sequences under supervised training and allows dictionary updates without requiring retraining.
-Liu [16] introduced the TFM network, aimed at enhancing character understanding by integrating positional information, thereby capturing both relative and absolute positional relationships among characters within the model to improve character interrelation comprehension.
-Besides, many studies also explore Chinese word structure, phonetics, and radicals as features for named entity recognition.
-Mai [17] examines different practical features of homophones based on lexical and phonetic characteristics.
-Li [18] leverages syntactic dependency trees to guide global attention mechanisms through rule-based methods.
-3 Methodology
-3.1 NER Task Definition
-The NER task is a typical sequence labeling task.
-Based on a $token$ sequence $x={x_{1},x_{2},\cdots,x_{n}}$ of a given length $n$, the output $y={y_{1},y_{2},\cdots,y_{n}}$ is predicted.
-Where $x_{i}\in\text{T}$, $T$ represents the vocabulary.
-In this study, the BIO labeling method is used to label the category $y$.
-The entire task flow requires the construction of a text processing sequence, as shown in the Equation-1.
-$input$ =$=\{x_{1},x_{2},\cdots,x_{n}\}$ Tokenization (1)
-$vectoried$ =$=\{\text{i}_{1},\text{i}_{2},\cdots,\text{i}_{n}\}$ BERT Embedding
-$featured$ =$=\{\text{v}_{1},\text{v}_{2},\cdots,\text{v}_{n}\}$ Feature Extraction
-$cls$ =$=\{\text{f}^{1}_{1},\text{f}^{1}_{2},\cdots,\text{f}^{C}_{n}\}$ Segmentation Weight
-$y$ =$=\{c_{1},c_{2},\cdots,c_{n}\}$ Sequence Dependency
-where $C$ denotes number of classes.
-In detail, the whole process starts with the input text sequence, and the text is converted into a set of $token$ through word segmentation.
-Then, these $token$ are mapped to the vector space through the BERT embedding layer to generate an embedded representation of each $token$.
-Subsequently, the embedded representation is passed to the feature extraction module to obtain deeper contextual features.
-The classification layer calculates multi-category scores for the feature vectors at each position to obtain the weights corresponding to each category.
-Finally, CRF layer is used to further model the dependencies between sequences, and the argmax operation is used to select the label sequence with the highest probability to obtain the final prediction output.
-3.2 Overview of HREB-CRF Architecture
-Proposed model is shown in Figure-2.
+Chinese NER faces unique challenges due to tokenization-induced semantic discrepancies [7].
+Character-based segmentation outperforms word-based approaches [8], and while Word2Vec [9] struggles with polysemy, hybrid models like CWVF-BiLSTM-CRF improve accuracy by combining character and word inputs [10].
+Adversarial transfer learning aids boundary detection [11], while Lattice-LSTM [1] and its variants (e.g., LR-CNN [12]) leverage dictionaries and attention to resolve lexical conflicts.
+Multi-representation encoding [13] and FLAT-Transformer’s lattice structures [2] further advance performance.
+Specialized data structures enhance NER feature capture.
+Tree-LSTM [14] integrates sub-features via topological networks, while DyLex [15] augments BERT with updatable dictionaries.
+Liu et al.’s TFM [16] improves character understanding through positional encoding.
+Besides, other works on exploiting Chinese linguistic features.
+Mai et al.[17] studies homophones, and Li et al.
+[18] uses syntactic dependency trees for global attention.
+III Methodology
+III-A Overview of HREB-CRF Architecture
+Proposed model is shown in Figure 1.
 BERT structure model is used as the pre-training head, and output the start and end (B-Cls and last I-Cls structure annotation) features of the CRF-constrained entity.
@@
 The left figure shows the end-to-end model architecture from text input to entity classification output, and the right figure shows the data processing flow of the RHEMA architecture.
-3.3 Exponentially Moving Average
+III-B Exponentially Moving Average
 Moving average method is widely used in sequence analysis tasks, such as stock and anomaly detection.
 It is proposed to reduce the global impact of local outliers anomalies.
-Generally speaking, most of the usage of MA are: simple moving average(Equation-2), weighted moving average(Equation-3), cumulative moving average(Equation-4) and exponentially weighted moving average(Equation-5).
+Generally speaking, most of the usage of MA are: simple moving average in Equation 1, weighted moving average in Equation 2, cumulative moving average in Equation 3 and exponentially weighted moving average in Equation 4.
 $SMA_{t}=\frac{1}{N}\sum_{i=t-N+1}^{t}x_{i}$ (1)
@@
 A larger $\alpha$ value indicates that current data is given more importance and the reliance on earlier data is weakened.
-Therefore, EM can effectively reduce the impact of random factors in the sequence on the overall trend, because as time goes by, $(1-\alpha)^{t}\cdot y_{t}$ will gradually decrease and tend to zero, gradually diluting the weight of earlier data.
+Therefore, EM can effectively reduce the impact of random factors in the sequence on the overall trend, because as shift window goes by, $(1-\alpha)^{t}\cdot y_{t}$ will gradually decrease and tend to zero, gradually diluting the weight of earlier data.
 The advantage of EM lies in its high responsiveness to recent data, which can adapt to data trend changes more quickly while reducing lag.
@@
 This feature gives EM sensitivity to changes in data trends, can smooth short-term fluctuations in the sequence, and retain information about longer trends, thereby playing a role in smoothing and denoising in the analysis and prediction of time series data.
-3.4 HEMA
+III-C HEMA
 Since the exponential averaging method is unique in the decay weight mechanism, we add the reduced-biased exponential averaging layer to the attention mechanism.
 Based on the MEGA method of [5], reduced-biased Hierarchical EMA is introduced as improved model.
-This improved architecture is shown in the Figure-3.
+This improved architecture is shown in the Figure 2.
 Figure 2: Proposed RHEMA architecture with added reduced-biased module.
 The figure illustrates the architecture of the RHEMA framework for text processing, enhanced by the addition of a reduced-bias module, marked in red.
-The input text is initially transformed through a token embedding layer.
 The core RHEMA block includes sequential processing through Batch Normalization and a FeedForward network.
-A novel "Reduced-Bias Add" module, shown with red arrows, is integrated to mitigate potential bias in the output, while preserving a feedback loop to earlier stages.
-This enhancement aims to improve robustness and fairness in predictions.
-The final output weights are generated post these computations.
+A novel ”Reduced-Bias Add” module, shown with red arrows, is integrated to mitigate potential bias in the output, while preserving a feedback loop to earlier stages.
 RHEMA uses GAU[19] to share feature dimensions.
-First, pass it to the EMA layer at Equation-6, then use the SiLU self-gated activation function to linearly project to z dimension, and add nonlinear features.
+First, pass it to the EMA layer at Equation 5, then use the SiLU self-gated activation function to linearly project to z dimension, and add nonlinear features.
 $X_{EMA}=EMA(X_{in})$ (5)
@@
 With the advantage of SiLU, the operation here can compress the originally redundant d-dim input into a more compact z-dim representation while retaining the context information.
-Equation-8 represents the specific calculation process of SiLU.
+Equation 7 represents the specific calculation process of SiLU.
 $F_{SiLU}=\sigma(x)+x\cdot\sigma(x)\cdot(1-\sigma(x))$ (7)
-Then, through the output of Equation-7, Query(Equation-9), Key(Equation-10) and Value(Equation-11) is calculated.
+Then, through the output of Equation 6, Query in Equation 8, Key in Equation 9 and Value in Equation 10 is calculated.
 Among them, the bias terms ($\mu_{q}$, $\mu_{k}$ and $b_{v}$) add additional degrees of freedom to the attention mechanism, which allows the query and key to adjust their respective baseline values based on the shared representation $Z$ to make them more suitable for broader tasks.
@@
 $V=F_{SiLU}(X_{in}W_{b}+b_{v})$ (10)
-Next, the result of the key-value pair query can be expressed as Equation-12.
+Next, the result of the key-value pair query can be expressed as Equation 11.
 $O=f_{S}(\frac{QK^{T}}{\daleth}+b_{rel})V$ (11)
 where $f_{S}(x)$ means single headed attention, and the bias term of $b_{rel}$ can be optimally solved by attention. $\daleth$ represents the scaling factor, which reduces the digits of large queries.
-by simulating the gating mechanism of GRU, activation output(Equation-13) is introduced. It affects the influence of each cell on the global affection.
+by simulating the gating mechanism of GRU, activation output in Equation 12 is introduced. It affects the influence of each cell on the global affection.
 $\hat{Y}=F_{SiLU}(X_{SiLU}W_{h}+(\gamma\odot O)U_{h}+b_{h})$ (12)
 Where the reset gate $\gamma$ controls whether the weight matrix $U_{h}$ affects the final output.
-Finally, the output is represented as Equation-14.
+Finally, the output is represented as Equation 13.
 $Y=\Phi\odot\hat{Y_{s}}+(1-\Phi)\cdot X_{in}$ (13)
 where $\Phi$ represents the update gate, smoothing the result by Sigmoid activation function.
-The single-head attention calculation method mentioned in Equation-14 can be expressed as Figure-4.
+The single-head attention calculation method mentioned in Equation 13 can be expressed as Figure 3.
 Figure 3: The architecture of single-head attention.
 The output $Y$ combines the candidate activation state $\hat{Y}$ and the original input $X_{in}$, and the update gate $\Phi$ controls the information fusion ratio between the two parameters.
-When calculating the candidate activation state $\hat{H}$, the exponential average attention mechanism uses the Laplace attention function(Equation-15) as the core component.
-Based on the concept of reduced-bias, it can be improved the reduced-bias Laplace function as Equation-16 to adapt to the introduction of deviation inside.
+When calculating the candidate activation state $\hat{H}$, the exponential average attention mechanism uses the Laplace attention function in Equation 14 as the core component.
+Based on the concept of reduced-bias, it can be improved the reduced-bias Laplace function as Equation 15 to adapt to the introduction of deviation inside.
 $F_{la}(x;\mu,\sigma)=\frac{1+\text{erf}\left(\frac{x-\mu}{\sigma\sqrt{2}}\right)}{2}$ (14)
@@
 This smoothing range is mapped to (0, 1) using the erf function, which improves the global context modeling ability of the model, and ultimately improves the expression effect and task performance of $Y$.
-3.5 Reduced-bias Module
+III-D Reduced-bias Module
 Compared with the traditional residual connection layer, proposed work have improved on the basis of the classic residual network structure (Residual Network)[20], and added a dynamic adjustment mechanism for the network weights before and after the residual connection.
@@
 In addition, this improvement can effectively alleviate the information transfer bias problem caused by fixed weights in traditional residual connections, and provide stronger robustness and flexibility for deep feature extraction and learning of the model.
-Classic resnet can be expressed as Equation-17.
+Classic resnet can be expressed as Equation 16.
 $\mathbf{y}_{l}=\mathcal{F}(\mathbf{x}_{l},\mathbf{W}_{l})+\mathbf{x}_{l}$ (16)
 where $\mathbf{x}_{l}$ is the input of the $l$-th layer module. $\mathcal{F}(\cdot)$ represents the operation of the network in this layer. $\mathbf{W}_{l}$ is a parameter of the $l$-th module.
-Among Equation-17, dynamic adaptation mechanism parameters $\alpha$ and $\beta$ is introduced. The new network can be expressed as Equation-18.
+Among Equation 16, dynamic adaptation mechanism parameters $\alpha$ and $\beta$ is introduced. The new network can be expressed as Equation 17.
 $\mathbf{y}_{l}=\alpha_{l}\cdot\mathcal{F}(\mathbf{x}_{l},\mathbf{W}_{l})+\beta_{l}\cdot\mathbf{x}_{l}$ (17)
@@
 where $\mathcal{L}$ is the loss function, $\mathbf{W}_{\alpha}$ and $\mathbf{W}_{\beta}$ are learnable weight matrices, $b_{\alpha}$ and $b_{\beta}$ are bias terms, and $\sigma(\cdot)$ is the activation function.
-By deriving Equation-19, $\mathbf{y}_{l}$ can be got by Equation-20.
+By deriving Equation 18, $\mathbf{y}_{l}$ can be got by Equation 19.
 $\mathbf{y}_{l}=$ $\,\sigma\left(\frac{\partial\mathcal{L}}{\partial\mathcal{F}(\mathbf{x}_{l},\mathbf{W}_{l})}\cdot\mathbf{W}_{\alpha}+b_{\alpha}\right)\cdot\mathcal{F}(\mathbf{x}_{l},\mathbf{W}_{l})$ (19)
@@
 Using optimization of the parameters of this layer, the contribution of different paths in the residual connection to the final output can be dynamically adjusted, thus achieving adaptive learning and weight optimization.
-3.6 Robustly Optimized BERT
+III-E Robustly Optimized BERT
 RoBERTa (Robustly Optimized BERT Approach)[6] is an improved version of BERT proposed by Facebook AI.
@@
 Compared with BERT, RoBERTa’s more powerful expression ability and flexible training strategy make it a pre-trained language model with outstanding performance in many downstream tasks.
-4 Experiments
-4.1 Datasets Description
+IV Experiments
+IV-A Datasets Description
 Three NER benchmark datasets are collected to verify the performance of proposed model. They are from Weibo[21], MSRA[22], and Resume[1].
-The specific properties of the original datasets are shown in the Table-1.
+The specific properties of the original datasets are shown in the Table I.
 It should be noted that since MSRA does not have a validation set by default, all experiments directly use the test-set as the valid-set.
-For the other two datasets, a split rate of 7:1.5:1.5 is used to evaluate.
-Table 1: Properties of the three datasets used.
+For the other two datasets, a split rate of 70%:15%:15% is used to evaluate.
+TABLE I: Properties of the three datasets used.
 Datasets Weibo MSRA Resume
 Class 8 3 8
-train 1350 46364 3821
-test 270 4365 477
-valid 269 4365 463
+Train 1,350 46,364 3,821
+Test 270 4,365 477
+Valid 269 4,365 463
 Avg length 54.61 46.80 32.47
@@
 Min length 7 5 3
-4.2 Experiment Settings
+IV-B Experiment Settings
 To verify the versatility and effectiveness of the proposed model in the NER task, BiLSTM-CRF is used BiLSTM-CRF as the baseline model.
@@
 The number of training epochs is set to 100 times with early stop callback.
-4.3 Evaluation Metrics
+IV-C Evaluation Metrics
 Proposed work uses precision (P), Recall (R), and F-Score ($F_{1}$) to evaluate the model.
-The negative log-likelihood function, shown as Equation-21 is considered as the loss function to quantitatively evaluate the performance of the model.
+The negative log-likelihood function, shown as Equation 20 is considered as the loss function to quantitatively evaluate the performance of the model.
 $\text{NLL}=-\sum_{i=1}^{N}\sum_{c=1}^{C}y_{i,c}\log(\hat{y}_{i,c})$ (20)
 where $y_{i,c}$ is the binary indicator (0 or 1) of whether class label $c$ is the correct classification for sample $i$, and $\hat{y}_{i,c}$ is the predicted probability of sample $i$ belonging to class $c$.
-4.4 Experimental Results
-4.4.1 Comparison Studies
-The experimental results on MSRA, Resume and Weibo for comparison studies are shown in the several tables respectively.
-Table 2: Comparison Experimental results on MSRA.
+IV-D Experimental Results
+IV-D1 Comparison Studies
+The experimental results on MSRA, Resume and Weibo for comparison studies are shown in the several tables.
+TABLE II: Comparison Experimental results on MSRA.
 Models Description MSRA
@@
 HREB (Ours) 94.58 93.98 94.28
-Compared with multiple existing methods on MSRA dataset by Table-2, he classic Lattice LSTM model achieves an F1 score of 93.18, demonstrating its effectiveness by combining character and word-level information.
+Compared with multiple existing methods on MSRA dataset by Table II, the classic Lattice LSTM model achieves an F1 score of 93.18, demonstrating its effectiveness by combining character and word-level information.
 However, proposed method surpasses this benchmark with an F1 score of 94.28, demonstrating stronger semantic modeling capabilities.
@@
 Comprehensive results show that proposed method achieves state-of-the-art performance on MSRA dataset.
-Table 3: Comparison Experimental results at Resume.
+TABLE III: Comparison Experimental results at Resume.
 Models Description Resume
@@
 HREB (Ours) 95.53 96.61 96.07
-Although proposed model has achieved meaningful recall (96.61) and F1 score (96.07) on the Resume dataset, detailed in Table-3.
+Although proposed model has achieved meaningful recall (96.61) and F1 score (96.07) on the Resume dataset, detailed in Table III.
 There are still some shortcomings that deserve attention.
@@
 96.92), indicating that there is space for improvement in the prediction accuracy of specific entities.
-Table 4: Comparison Experimental results at Weibo.
+TABLE IV: Comparison Experimental results at Weibo.
 Models Description Weibo
@@
 Glyce Meng 2019 [30] 67.68 67.71 67.60
-FLAT Li 2020[2] 63.42
-FLAT+BERT Li 2020[2] 68.55
+FLAT Li 2020 [2] - - 63.42
+FLAT+BERT Li 2020 [2] - - 68.55
 MSFM Zhang 2022 [26] 60.75 51.83 55.94
@@
 Figure 4: Based on the web search results, a piece of news was randomly selected as a test sample, which contains the extraction of names of people, places and organizations.
-Proposed method achieves an F1 score of 68.60, slightly higher than FLAT+BERT (68.55) and close to Glyce (67.60), demonstrating robustness in complex text processing on Weibo, shown as Table-4.
+Proposed method achieves an F1 score of 68.60, slightly higher than FLAT+BERT (68.55) and close to Glyce (67.60), demonstrating robustness in complex text processing on Weibo, shown as Table IV.
 Furthermore, proposed method performs well in recall (70.76), indicating its advantage in identifying more potential entities.
 However, compared with the precision of DAE (69.68), proposed method still has a certain gap in accuracy, which suggests that proposed model still has room for improvement in accurate prediction when dealing with noisy social media data.
-4.4.2 Ablation Studies
-In the ablation experiment, shown as Table-5, two experiments are designed to evaluate the real impacts of proposed structure.
-First, each module is removed to confirm the synergy between different modules.
-Second, BERT is replaced or removed to compare the impact of word embedding weights on the next layer input.
+IV-D2 Ablation Studies
+Ablation Study in Table V evaluates the proposed structure through two experiments: (1) module removal to assess synergy, and (2) BERT replacement/removal to analyze word embedding impacts.
 Comparison between the base model and different BERT models
 The experimental results of comparing the base model (S.N.1) with the models that introduce different BERT variants (S.N.2 and S.N.3) show that RoBERTa outperforms BERT in terms of vocabulary vectorization.
-This shows that RoBERTa has a stronger ability to represent text semantics and can capture contextual information more effectively.
 Comparison between naive attention and hierarchical EMA
@@
 In the comparison of the final model (S.N.8) with the ordinary BERT model (S.N.7), the experimental results show that even with the introduction of a rich feature extraction module, RoBERTa can still efficiently capture the deep features of the text and show a superior feature learning ability.
-Table 5: Ablation Studies on MSRA
+TABLE V: Ablation Studies on MSRA
 S.N. Module Metric
@@
 7 $\surd$ $\surd$ BERT-base 93.55 93.88 93.71
-8
-$\surd$ $\surd$ RoBERTa 94.58 93.98 94.28
-4.5 Parameter Selection
+8 (Ours) $\surd$ $\surd$ RoBERTa 94.58 93.98 94.28
+IV-E Parameter Selection
 In the design of the attention module based on the EMA (Exponential Moving Average) mechanism, the number of attention heads is considered to be one of the important factors affecting the model performance.
@@
 The experimental results show that when the number of attention heads is set to 64, the model achieves the best performance in performance indicators such as precision, recall, and F1 score.
-Table 6: Effects of different parameters on performance in RHEMA.
+TABLE VI: Effects of different parameters on performance in RHEMA.
 EMA head Metric Dataset
@@
 F1 94.21 68.06 95.45
-As seen from the Table-6, too few attention heads with 16 or 32, may cause the model to be unable to fully capture feature representations, thereby limiting the performance of the model.
+As seen from the Table VI, too few attention heads with 16 or 32, may cause the model to be unable to fully capture feature representations, thereby limiting the performance of the model.
 When the number of attention heads increases to 64, the performance of the model reaches its peak.
@@
 This phenomenon may be due to the fact that too many attention heads increase the complexity of the model, thereby introducing the risk of overfitting.
-4.6 Case Analysis
+IV-F Case Analysis
 Real Example
 Select text data on the Internet for analysis.
-In Example-1 of Figure-5, the performance of the model trained on MSRA dataset is verified using normal press releases.
+In Example-1 of Figure 4, the performance of the model trained on MSRA dataset is verified using normal press releases.
 The results show that proposed model can effectively extract entity information for regular text.
@@
 This case involves the extraction of obfuscated names.
-In Example-2 of Figure-7, an extreme scenario was constructed.
+In Example-2 of Figure 6, an extreme scenario was constructed.
 The same Chinese character was expressed as a noun, a verb, and an adjective simultaneously.
@@
 Select text data on the datasets for analysis.
-The analysis presented in Figure-6 addresses two primary challenges in entity recognition, as outlined in the introduction: 1) ambiguity in boundary identification (example-3); 2) errors in entity recognition (example-4).
+The analysis presented in Figure 5 addresses two primary challenges in entity recognition, as outlined in the introduction: 1) ambiguity in boundary identification (example-3); 2) errors in entity recognition (example-4).
 A comparative evaluation demonstrates that the proposed model effectively resolves these issues in the selected examples, exhibiting superior performance relative to the baseline model.
-4.7 Analysis and Discussion
-Through comparative experiments and ablation studies, an in-depth analysis of various aspects of model performance was conducted.
-The following are the main findings and discussions derived from the experimental results:
-1) Proposed method can be matched with the comparative experiment and the ablation experiment to prove that the model can adapt to different types of semantic information under different scene tasks and adaptively improve the robustness of CNER.
-Comparing the experimental results of Table-2, Table-4, and Table-3, proposed model achieves the best performance in both regression and F1 indicators.
-The ablation experiment in Table-5 shows that the hierarchical exponential attention and reduced-biased modules can effectively expand the deep dimensions of the data compared with the baseline model.
-It performs well in the conventional tasks in Figure-5.
-2) Proposed research shows that the Hierarchical Reduced-bias EMA architecture can bring accurate feature extraction for CNER tasks.
-The modeling of the attention mechanism is to understand the text content from the vector dimension, which is used in the latest LLMs models (such as GPT, Llama).
-In the field of transfer learning, inspired by DAE[27], proposed model can be used in other sequence modeling tasks.
-The sequence modeling task uses BERT to model the text vectorization input, and the output is the probability weight of the same length as the input text.
-It is believed that the model can be applied to NER tasks such as text understanding, text classification, and emotion recognition simultaneously.
-3) Increasing the number of attention heads will significantly increase the number of parameters in the multi-head attention mechanism, thereby increasing the training time and inference time of the model.
-When selecting the number of attention heads, proposed study comprehensively considered the trade-off between performance and efficiency and selected EMA head of 64.
-5 Conclusion
+IV-G Analysis and Discussion
+Through comparative and ablation studies, key findings on model performance are summarized:
+1) Proposed model adapts to diverse semantic information across scene tasks, enhancing CNER robustness.
+As shown in Tables II, IV and III, proposed method achieves top performance in regression and F1 scores.
+Ablation studies in Table V confirm that hierarchical exponential attention and reduced-bias modules expand data depth, excelling in conventional tasks in Figure 4 and domain tasks in Figures 6 and 5.
+2) The Hierarchical Reduced-bias EMA architecture enables precise CNER feature extraction.
+Inspired by DAE [27], the model extends to sequence tasks (e.g., text classification, emotion recognition) using BERT-based vectorization.
+3) Increasing attention heads raises parameters, prolonging training/inference time. In order to balanceto balance performance and efficiency, 64 EMA heads were selected.
+V Conclusion
 Proposed study focus on CNER model that combines RoBERTa, reduced-biased Hierarchical EMA, BiLSTM, and CRF, which conducts extensive experimental verification on multiple Chinese datasets.
@@
 Acknowledgments
-This was was supported in part by Agency for Science, Technology and Research.
-Appendix
-5.1 CRF Decoding
-CRF was introduced to strengthen the shortcomings of HMM in feature selection and dependency assumptions.
-HMM is a generative model that models the joint probability $P(x,y)$, that is, the joint distribution of the observed data and the label sequence, expressed as the Equation-22.
-This requires HMM to make strong assumptions about the relationship between the observed data and the hidden state (such as the independent and identically distributed assumption and the Markov assumption).
-In contrast, CRF is a discriminative model that directly models the conditional probability $P(y|x)$ without making independence assumptions on the observed data.
-Therefore, CRF introduces feature functions and weights into the model and maximizes the conditional probability of given observed data by learning these parameters.
-It is able to directly model the label sequence without assuming the Markov property between the observed data and the hidden state.
-This enables CRF to better capture long-distance dependencies in the sequence and automatically learn the relationship between features without manually designing features.
-$P(x,y)=P(y_{1})\prod_{t=2}^{T}P(y_{t}|y_{t-1})\prod_{t=1}^{T}P(x_{t}|y_{t})$ (22)
-The calculation process of CRF involves constructing and optimizing the conditional probability $P(y|x)$, that is, the probability of predicting the label sequence $y$ given the observation sequence $x$.
-The core of CRF is to maximize this conditional probability by defining feature functions and learning weights.
-Specifically, the derivation formula of CRF is:
-$P(y|x)=\frac{1}{Z(x)}\exp\left(\sum_{t=1}^{T}\sum_{k}\lambda_{k}f_{k}(y_{t-1},y_{t},x,t)\right)$ (23)
-where $f_{k}(y_{t-1},y_{t},x,t)$ is the feature function that defines the relationship between the label sequence $y$ and the observation sequence $x$.
-$\lambda_{k}$ is the weight parameter of the feature function, which is learned through training.
-$Z(X)$ is the normalization term (partition function) that ensures that the sum of the conditional probabilities is the Equation-24.
-$Z(x)=\sum_{y}\exp\left(\sum_{t=1}^{T}\sum_{k}\lambda_{k}f_{k}(y_{t-1},y_{t},x,t)\right)$ (24)
-5.2 Model Training Performance
-Figure-8, Figure-9 and Figure-10 are the training processes of MSRA, Weibo and Resume respectively.
-The training process shows that MSRA reaches asymptotic performance at about 60 steps, Weibo reaches asymptotic performance at 15 steps, and Resume reaches asymptotic performance at 85 steps.
-Figure 8: Training progress metrics for MSRA.
-Figure 9: Training progress metrics for Weibo.
-Figure 10: Training progress metrics for Resume.
+This research was funded by Programme MARS (Programme of Maritime AI Research in Singapore) with funding grant number SMI-2022-MTP-06 by Singapore Maritime Institute (SMI).